# I’m Sorry Dave, I’m Afraid I Can’t Return That: On YouTube Search API Use in Research

### v1

@@ 6 Discussion
In this short paper, we contribute to our understanding of a crucial research tool. The behaviors we record offer both new strategies of working with the YouTube API, as well as opening further research directions to better understand how researchers can use it.
6.1 Implications
@@
Our findings show that binning queries across the observation period is not as fruitful as previously thought [22, 24, 25], and offers low return on investment considering the quota cost of the search endpoint. Instead, researchers may experiment with breaking up their topics as opposed to their time frames. This can be achieved by incorporating more AND statements in the query or querying for multiple sub-topics (e.g., specific players alongside their national teams instead of the entirety of the World Cup event). The total number of results in the query metadata is a crucial way of assessing how optimal a query is (with lower being better/more stable).
Moreover, researchers who use the API for data collection should aim to use ID-based endpoints wherever possible, as these are more complete and reliable. The strategy of pre-selecting channels, provided that there are ways to obtain an adequately representative sample, is a viable one as long as the search endpoint is not used to collect their videos. Instead, as we outline in Section 2, researchers can use a combination of the Channels: list and PlaylistItems: list endpoints, both of which take IDs as queries, to get complete channel uploads.
Lastly, we urge researchers to clearly outline which endpoints they use in their data collection pipelines, as these choices can massively impact the replicability of their work.
6.2 Limitations and Future Work
The new strategies we put forward may need experimental validation, for example, by running progressively more restrictive queries and seeing how that influences the replicability of the data returned (alongside the reported video pool size).
Moreover, we do not explicitly analyze how other parameters, such as ordering or including channel IDs, may affect replicability, although we have no reason to believe that this would alter our conclusions (especially as reverse chronological order should be immutable for historical queries, other than video deletions).
@@ 6.2 Limitations and Future Work
Finally, there are several future directions that scholars can pursue to better understand this invaluable tool. For example, future research can replicate our experiments with more sparse collections over a longer period, to check for potential periodicity in set similarities. Moreover, given the substantial efforts that scholars have expended in creating sockpuppets for YouTube SERP audits [7, 8], future research can employ similar methods to ours to check the consistency between results of sockpuppet SERPs and search endpoint results. This would help us understand if the search endpoint has research value beyond data collection, for example, as a low-resource way of conducting SERP audits.
@@
Acknowledgements. This work has been supported by the University of Washington’s Center for an Informed Public, the John S. and James L. Knight Foundation (G-2019-58788), and the William and Flora Hewlett Foundation (2023-02789). The author is grateful to Adrian Padilla, Bernhard Rieder, and Ethan Zuckerman for feedback, comments, and discussions related to this work.
Appendix A Query Parameters
@@ A.1 General parameters
A.2 Topic-specific parameters
@@
Appendix B ID-Based Queries
@@
B.1 Video: list Endpoint
@@
Figure 4: Parallel plots of percentage of common videos retrieved between successive runs and overall Jaccard similarity of common videos retrieved using the “Videos: list” endpoint. Comparison ID refers to the respective collection (higher ID = later collection).
B.2 CommentThreads: list and Comments: list Endpoints
@@
Table 5: Jaccard similarities between first- and last-collection comment sets. TL = top-level, N = nested. NS = non-shared videos (full sets), S = shared videos only. N/A values for the Higgs topic, which is the oldest, are possibly due to the comment reply affordance operating differently in 2012.
Appendix C Regression Robustness Checks
Alternative regression model setups as robustness checks against the binned regression model implemented in the main paper.
C.1 Frequency as Continuous Variable
We use frequency as our dependent variable in a multiple Ordinary Least Squares (OLS) regression with robust standard errors. The overall model is significant ($F_{(14,5348)}=122.3,p<0.001$) and shows modest fit ($R^{2}=0.164$). We report standardized beta coefficients with confidence metrics in Table 6. The patterns are identical to what is reported in the main paper.
@@ C.1 Frequency as Continuous Variable
Table 6: Standardized regression coefficients for OLS model. *$p<0.05$, **$p<0.01$, ***$p<0.001$.
C.2 Non-Binned Ordinal Regression
@@
Table 7: Standardized regression coefficients for full ordered model. *$p<0.05$, **$p<0.01$, ***$p<0.001$.

### v2

@@ 6. Discussion
In this short paper, we contribute to the understanding of a crucial research tool. The behaviors we record offer new strategies of working with the YouTube API and future research directions on expanding its use-cases.
6.1. Implications
@@
Our findings show that binning queries across the observation period is not as fruitful as previously thought (Rieder et al., 2018; Violot et al., 2024; Yin and Brown, 2018), and offers low return on investment considering the quota cost of the search endpoint. Instead, researchers may experiment with breaking up their topics as opposed to their time frames. This can be achieved by incorporating more AND statements in the query or querying for multiple sub-topics (e.g., specific players alongside their national teams instead of the entirety of the World Cup event). The total number of results in the query metadata is a crucial way of assessing how optimal a query is (with lower being better/more stable). To reduce this number of results as much as possible, researchers should build their queries around API parameters that are specified in the documentation to affect search results (e.g., query, region code, etc.), and not merely the API response (e.g., before-after datetimes, topic IDs, etc.)
Alternatively, in cases where discovery is possible at the channel level instead of using keywords, or where consistency takes precedence over data completeness, ID-based endpoints are a viable route. For example, as we outline in Section 2, a combination of the Channels: list and PlaylistItems: list endpoints, both of which take IDs as queries, would allow for the retrieval of complete channel uploads. We also urge researchers to specify these endpoints in their data collection pipelines, as these choices can massively impact the replicability of their work.
6.2. Limitations and Future Work
The strategy of using parameters that restrict the potential results pool is not explicitly quantified here, and can be addressed in future work that designs progressively restrictive queries using other parameters as we do here with time constraints. Moreover, we do not explicitly analyze how other parameters, such as ordering or including channel IDs, may affect replicability.
@@ 6.2. Limitations and Future Work
Future work can also replicate our experiments with more sparse collections over a longer period, to check for potential periodicity in set similarities. Moreover, given the substantial efforts that scholars have expended in creating sockpuppets for YouTube SERP audits (Hussein et al., 2020; Jung et al., 2025), similar methods to ours can be employed to check the consistency between results of sockpuppet SERPs and search endpoint results. This would help us understand if the search endpoint has research value beyond data collection, for example, as a low-resource way of conducting SERP audits.
Acknowledgments
@@
This work has been supported by the University of Washington’s Center for an Informed Public, the John S. and James L. Knight Foundation (G-2019-58788), and the William and Flora Hewlett Foundation (2023-02789). The author is grateful to Adrian Padilla, Bernhard Rieder, Ethan Zuckerman, and the anonymous reviewers for their helpful feedback.
Appendix A Ethics
This work makes sole use of video, comment, or channel IDs and high-level metadata (number of views, likes, etc.) as data points and does not analyze content beyond this point. The aim of this paper is to offer a better understanding of how data from YouTube, one of the largest platforms worldwide, should be understood and used by researchers. The societal benefits arising from this work apply insofar as the important topics studied on YouTube that we cover in the Introduction benefit from the directions we offer.
Appendix B Query Parameters
@@ B.1. General parameters
B.2. Topic-specific parameters
@@
Appendix C ID-Based Queries
@@
C.1. Video: list Endpoint
@@
Figure 5. Parallel plots of percentage of common videos retrieved between successive runs and overall Jaccard similarity of common videos retrieved using the “Videos: list” endpoint. Comparison ID refers to the respective collection (higher ID = later collection).
C.2. CommentThreads: list and Comments: list Endpoints
@@
Table 4. Jaccard similarities between first- and last-collection comment sets. TL = top-level, N = nested. NS = non-shared videos (full sets), S = shared videos only. N/A values for the Higgs topic, which is the oldest, are possibly due to the comment reply affordance operating differently in 2012.
Appendix D Regression Robustness Checks
Alternative regression model setups as robustness checks against the SHAP-explained gradient boosting model implemented in the main paper.
D.1. Binned Ordinal Regression
We split the frequency of returns into four roughly equal bins (1-5, 6-10, 11-15, 16), taking into account that 16 is the modal value. We then perform an ordinal regression using a logit link function. This function is chosen due to the uniform distribution arising from the binning. All continuous features are log-transformed to reduce multicollinearity and standardized for better comparison between coefficients.
Variable $\beta$ SE 95% CI
SD (quality) -0.018 0.079 [-0.173, 0.137]
brexit (topic) ***1.231 0.098 [1.039, 1.423]
capriot (topic) -0.160 0.093 [-0.341, 0.022]
grammys (topic) *0.171 0.083 [0.008, 0.333]
higgs (topic) ***3.10 0.141 [2.826, 3.379]
worldcup (topic) 0.161 0.101 [-0.037, 0.359]
duration ***-0.115 0.028 [-0.170, -0.061]
views 0.161 0.088 [-0.011, 0.333]
likes **0.285 0.095 [0.098, 0.471]
comments 0.069 0.064 [-0.058, 0.195]
channel age 0.049 0.031 [-0.012, 0.110]
channel views *0.3176 0.135 [0.053, 0.582]
channel subs **-0.3784 0.122 [-0.617, -0.140]
# channel videos -0.0212 0.075 [-0.169, 0.126]
Table 5. Standardized regression coefficients for binned ordinal model. *$p<0.05$, **$p<0.01$, ***$p<0.001$.
D.2. Frequency as Continuous Variable
@@
Table 6. Standardized regression coefficients for OLS model. p<0.05*p<0.05, p<0.01**p<0.01, p<0.001***p<0.001.
We use frequency as our dependent variable in a multiple Ordinary Least Squares (OLS) regression with robust standard errors. The overall model is significant ($F_{(14,5348)}=122.3,p<0.001$) and shows modest fit ($R^{2}=0.164$). We report standardized beta coefficients with confidence metrics in Table 6. The patterns are identical to the previous binned ordinal regression model and to what is reported in the main paper.
D.3. Non-Binned Ordinal Regression
@@
Table 7. Standardized regression coefficients for full ordered model. *$p<0.05$, **$p<0.01$, ***$p<0.001$.